# Covariate balancing estimation and model selection for difference-in-differences approach

### v1

@@ 5.1 Derivation of Model Selection Criterion
As in the case with known propensity scores in Section 4.2, we derive the model selection criteria when using the CBD method. We evaluate
@@
which is the third term of the risk function (17). First, substituting $\bm{x}_{i}^{{\rm T}}\bm{\theta}^{*}$ for ${\rm E}[y^{[1]}_{i}-y^{[0]}_{i}\mid\bm{x}_{i},d^{[1]}_{i}=1]$, and then substituting the right-hand side of (15) for $\hat{\bm{\theta}}^{\rm CBD}-\bm{\theta}^{*}$, this third term can be written as
$2{\rm E}\bigg{[}\frac{1}{\sqrt{n}}\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})\{\rho(\bm{t}_{i},\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})\Delta_{i}-\bm{x}_{i}^{{\rm T}}\bm{\theta}^{*}\}\bm{x}_{i}^{{\rm T}}$
@@
Since $(1/n)\sum_{j=1}^{n}e^{[1]}(\bm{x}_{j};\hat{\bm{\alpha}}^{\rm CB})\bm{x}_{j}\bm{x}_{j}^{{\rm T}}$ converges to
$\bm{L}(\bm{\alpha}^{\dagger})\equiv{\rm E}\{e^{[1]}(\bm{x};\bm{\alpha}^{\dagger})\bm{x}\bm{x}^{{\rm T}}\}$
in probability, if we let $\bm{u}$ be the normal random vector that is the weak limit of
@@
then the weak limit of (17) removing the expectation is expressed as
@@
In the setting where the estimator $\hat{\bm{\theta}}^{\rm CBD}$ is consistent with $\bm{\theta}^{*}$, ${\rm E}[\bm{u}\bm{u}^{{\rm T}}]={\rm Var}[\bm{u}]$ holds because the expected value of $\bm{u}$ becomes $\bm{0}$, and the asymptotic evaluation of (17), which is to be used in the information criterion, is expressed as
@@
Here, we write down ${\rm Var}[\bm{u}]$. As the error of the parameter $\hat{\bm{\alpha}}^{\rm CB}$ in the propensity score can be shown to be (14), using this, (21) can be expanded as
@@
$\ \ \ \ \ +\dfrac{\partial}{\partial\bm{\alpha}^{{\rm T}}}\bigg{(}e^{[1]}(\bm{x}_{i};\bm{\alpha}^{\dagger})\bigg{[}\frac{d^{[1]}_{i}-e^{[1]}(\bm{x}_{i};\bm{\alpha}^{\dagger})}{e^{[1]}(\bm{x}_{i};\bm{\alpha}^{\dagger})\{1-e^{[1]}(\bm{x}_{i};\bm{\alpha}^{\dagger})\}}\Delta_{i}-\bm{x}_{i}^{{\rm T}}\bm{\theta}^{*}\bigg{]}\bigg{)}(\hat{\bm{\alpha}}^{\rm CB}-\bm{\alpha}^{\dagger})\bigg{\}}\{1+{\rm o}_{\rm P}(1)\}$
$=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\Big{[}e^{[1]}(\bm{x}_{i};\bm{\alpha}^{\dagger})\{\rho(\bm{d}_{i},\bm{x}_{i};\bm{\alpha}^{\dagger})\Delta_{i}-\bm{x}_{i}^{{\rm T}}\bm{\theta}^{*}\}\bm{x}_{i}$
$\ \ \ \ \ \ \ \ \ \ -\bm{M}(\bm{\alpha}^{\dagger},\bm{\theta}^{*})\{\bm{G}^{{\rm T}}(\bm{\alpha}^{\dagger})\bm{W}\bm{G}(\bm{\alpha}^{\dagger})\}^{-1}\bm{G}^{{\rm T}}(\bm{\alpha}^{\dagger})\bm{W}\bm{h}(\bm{d}_{i},\bm{x}_{i};\bm{\alpha}^{\dagger})\Big{]}\{1+{\rm o}_{\rm P}(1)\},$
@@
it can be shown to be ${\rm Var}[\bm{u}]=\bm{V}(\bm{\alpha}^{\dagger},\bm{\theta}^{*})$. Thus, the following theorem is obtained.
@@
When the propensity score is estimated by the CBD method, the third term of (16) is evaluated as
${\rm E}[c^{\rm limit}]=2{\rm tr}\{\bm{L}(\bm{\alpha}^{\dagger})^{-1}\bm{V}(\bm{\alpha}^{\dagger},\bm{\theta}^{*})\}.$
As a result, when the propensity score is estimated by the CBD method, we propose
$\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})(\rho(\bm{d}_{i},\bm{x}_{i})(\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})\Delta_{i}-\bm{x}_{i}^{{\rm T}}\hat{\bm{\theta}}^{\rm CBD})^{2}+2{\rm tr}\{\bm{L}(\bm{\alpha}^{\dagger})^{-1}\bm{V}(\bm{\alpha}^{\dagger},\bm{\theta}^{*})\}$ (23)
as the model selection criterion for the SDID method, which is an asymptotically unbiased estimator of (16).
Also in this case, the penalty term depends on the true values of the parameters, thus empirical estimates are used in practice. Specifically, $\bm{L}_{n}(\hat{\bm{\alpha}}^{\rm CB})\equiv n^{-1}\sum_{i=1}^{n}\allowbreak e^{[1]}(\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})\bm{x}_{i}\bm{x}_{i}^{{\rm T}}$ is used instead of $\bm{L}(\bm{\alpha}^{\dagger})$ and
$\bm{V}_{n}(\hat{\bm{\alpha}}^{\rm CB},\hat{\bm{\theta}}^{\rm CBD})\equiv\frac{1}{n}$ $\sum_{i=1}^{n}\Bigl{[}e^{[1]}(\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})\{\rho(\bm{d}_{i},\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})\Delta_{i}-\bm{x}_{i}^{{\rm T}}\hat{\bm{\theta}}^{\rm CBD}\}\bm{x}_{i}$
$-\bm{M}_{n}(\hat{\bm{\alpha}}^{\rm CB},\hat{\bm{\theta}}^{\rm CBD})\{\bm{G}_{n}^{{\rm T}}(\hat{\bm{\alpha}}^{\rm CB})\bm{W}_{n}\bm{G}_{n}(\hat{\bm{\alpha}}^{\rm CB})\}^{-1}\bm{G}_{n}^{{\rm T}}(\hat{\bm{\alpha}}^{\rm CB})\bm{W}_{n}\bm{h}(\bm{d}_{i},\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})\Bigr{]}^{2}$
is used instead of $\bm{V}(\bm{\alpha}^{\dagger},\bm{\theta}^{*})$, where
$\bm{M}_{n}(\bm{\alpha},\bm{\theta})\equiv\frac{1}{n}\sum_{i=1}^{n}\dfrac{\partial}{\partial\bm{\alpha}}e^{[1]}(\bm{x}_{i};\bm{\alpha})\biggl{(}\frac{d^{[1]}_{i}-1}{(1-e^{[1]}(\bm{x}_{i};\bm{\alpha}))^{2}}\Delta_{i}-\bm{x_{i}}^{{\rm T}}\bm{\theta}\biggr{)}\bm{x_{i}}.$
Here, the propensity score is estimated using the covariate balancing approach, but the information criterion can be derived in the same way when the MLE is used (see Supplementary material).
@@ 5.2 Numerical experiments
Through simulations, we examine the accuracy of the approximation of the penalty term and the performance of the model selection criterion (23) based on Theorem 3 for the proposed CBD method. As the weighting matrix in the generalized method of moments, we use the identity matrix based on the results of Section 4.4. But even when an optimal matrix is used as the weighting matrix, the accuracy of the approximation of the penalty term and the performance of the model selection are comparable to those obtained when the identity matrix is used (see Supplementary Material).
In Table 4, data are generated as in Case 1-1 and Case 1-2 of Section 4.4 to verify whether the second term of (23), the penalty term of the proposed criterion, accurately approximates the bias (17) that we originally wanted to evaluate for some parameter values $\beta^{*}$ and sample size $n$. In all cases, the penalty term of the proposed criterion provides values much closer to the bias than the penalty term of the QICw, and its approximation accuracy is confirmed to be high. Conversely, in all cases, the penalty term of the QICw significantly underestimates the bias.
Table 4: Bias evaluation when the propensity score is estimated by the CBD method with identity matrix as the weighting matrix: True is the true value of the bias evaluated by the Monte Carlo method, and Proposal is the penalty term of the proposed method. The penalty term of the QICw is also given as a reference.
@@
In Table 5, data are generated as in Case 2-1, Case 2-2 and Case 2-3 of Section 4.4 and the proposed criterion and the QICw are compared by model selection using the forward selection method. The main index of the goodness of the model selection criteria is again the empirical estimate of the risk. Unlike the results in Section 4.4, where the propensity score is known, the proposed criterion always outperforms the QICw, even in Case 2-2, where relatively few covariates are unnecessary among the candidates. On the other hand, as in Case 2-1 and Case 2-3, the proposed criterion is particularly superior when the candidates include many covariates that do not contribute to the true structure, which is similar to Section 4.4. The penalty term is underestimated in the QICw, therefore, too many extraneous variables must be selected.
Table 5: Comparison of the proposed criterion and the QICw through model selection when the propensity score is estimated using the CBD method with identity matrix as the weighting matrix: Risk is the empirical estimate of the risk of the selected model, TP is the average number of covariates that are true positives, and FP is the average number of covariates that are false positives.
@@ 6 Real data analysis
The LaLonde dataset is treated in LaLonde (1986) and included in the R package Matching version 4.10.14 (Sekhon 2011). The group that took the U.S. job training program in 1976 is denoted as $t=1$ and the group that did not take the program is denoted as $t=0$. The difference in annual income in 1978 after the training for each group is estimated via the ATT. Age (age, years of education (educ), income in 1974 (re74), black (black), hispanic (hisp), married (married), and high school graduate or higher (nodegr) are treated as covariates, i.e. $p=7$, and the outcome variable is the difference between income in 1978 (re78) and income in 1974 (re78). The sample size is $N=445$. The objective of this section is not to extract information from this data, but to examine whether the proposed method makes a significant difference from existing methods, and therefore, a sufficiently small number of samples are assumed. Specifically, for a natural number $k$, the $23k-$ 2 rows of data are analyzed in Block 1, $13k-$ 1 rows in Block 2, and $3k$ rows in Block 3, respectively.
Regarding the estimation, we first estimate the propensity score using the CBD method with the identity matrix as the weighting matrix. Then, assuming that the ATT can be written as a linear sum of covariates, we estimate its coefficients by the SDID method. Table 5 selects covariates using the proposed criterion or the QICw with the forward selection method and estimates the coefficients of the selected variables. The QICw selects all variables in any Block, while the proposed criterion does not select some variables. In particular, the difference between the models selected for Block 1 and Block 3 is remarkable. Because we do not know the true structure of the real data, it is impossible to judge the superiority or inferiority of the two criteria, however, we can see that the difference between them is quite large. In a setting such as this article, it is essential to use model selection criteria that are theoretically valid rather than intuitive ones.
Table 6: Comparison of the proposed criterion and the QICw when propensity scores are estimated using the CBD method with the identity matrix as the weighing matrix, through model selection experiments on LaLonde data. The numbers with exponential notation are the estimated coefficients for each variable, and a 00 means that the variable was not selected in the model selection.
@@
QICw 2.38E+4 --1.91E+2 8.22E+2 2.73E-1
3 Proposal --8.50E+2 0E+0 0E+0 0E+0
QICw --8.51E+3 6.13E+0 3.35E+2 2.56E-1
@@ 1 Proposal 0E+0 0E+0 0E+0 0E+0
QICw 3.55E+2 --8.19E+3 --9.41E+2 --1.30E+3
2 Proposal --1.86E+4 --2.34E+4 4.02E+3 --1.26E+4
QICw --1.85E+4 --2.42E+4 4.88E+3 --1.34E+4
3 Proposal 0E+0 --1.03E+4 0E+0 0E+0
QICw 3.84E+3 --6.71E+3 2.92E+3 --5.90E+2
@@ 7 Discussion
In this article, we have proposed the CBD method, a SDID method incorporating covariate balancing. In order to ensure the doubly robustness, i.e., that the ATT can be consistently estimated if the model for the propensity score is correctly specified or the model for the changes in the outcomes over time is correctly assumed, we have shown that it is sufficient to balance the distribution up to the second-order moment of the covariates across groups with the treatment group as the target population. This fact is important to note because covariate balancing usually balances first-order moment. We also have developed model selection criteria for the CBD method as an asymptotically unbiased estimator of the risk based on the loss function used for the estimation. Specifically, we have derived model selection criteria using a DID-specific assumption called the conditional parallel trend assumption, i.e., without relying on the ignorable treatment assignment condition, which is often assumed in causal inference, and we have shown that the penalty term is considerably different from twice the number of parameters that often appears in AIC-type model selection criteria. Since there have been no model selection criteria for even the basic SDID method, we have developed the model selection criteria not only for the CBD method, but also for when the propensity score is known or the propensity score is estimated by the MLE. Numerical experiments showed that the proposed methods estimate ATT robustly compared to the method that estimates propensity scores by the MLE, and that the proposed criteria clearly reduce the risk targeted by the SDID method compared to the intuitive generalization of the existing information criterion to the SDID method. In addition, the real data analysis has confirmed that there were large differences between the results of the proposed method and the existing methods.
Although this article deals with a linear model of covariates as a simple working model for the outcome, it should be noted that the results can be easily extended to situations where a nonlinear model of covariates should be used, for example, when using the kernel method. In fact, Hazlett (2020) treats the function $\phi(\bm{x}):\mathbb{R}^{p}\rightarrow\mathbb{R}^{q}$ in the reproducing kernel Hilbert space as $q$ may be larger than $p$ or $n$. And they assume that the outcome is represented by a linear model of $\phi(\bm{x})$, the feature map of $\bm{x}$, which in our setting can be written as
${\rm E}[y^{[1]}(1)-y^{[0]}(1)\mid\bm{x},d^{[1]}=1]=\phi(\bm{x})^{{\rm T}}\bm{\theta}.$ (24)
@@
$\sum_{i=1}^{n}\bigg{\{}\frac{d^{[1]}_{i}\phi(\bm{x}_{i})}{e^{[1]}(\bm{x}_{i};\bm{\alpha})}-\frac{d^{[0]}_{i}\phi(\bm{x}_{i})}{e^{[0]}(\bm{x}_{i};\bm{\alpha})}\bigg{\}}=0$
as the moment condition for estimating the propensity score, they propose the kernel balancing method, which can flexibly estimate causal effects even if the relationship between the covariate and the outcome is nonlinear. In the setting of this article, we also consider the situation (24) and when making the estimation doubly robust, we assume that the change in the outcomes follows
${\rm E}[\Delta^{[k]}\mid\bm{x},d^{[k]}=1]=\phi(\bm{x})^{{\rm T}}\bm{\beta}^{[k]*}.$
Let $k(\cdot,\cdot):\mathbb{R}^{p}\times\mathbb{R}^{p}\rightarrow\mathbb{R}$ be the kernel function corresponding to $\phi(\bm{x})$, and let $\bm{K}$ be a $p\times p$ semidefinite symmetric matrix whose $(l,m)$ component is $k(x_{l},x_{m})$. If we estimate the parameter $\bm{\alpha}$ that satisfies
@@
and then estimate $\bm{\theta}$ or derive model selection criteria based on it, that would be a natural extension of the proposed method in this article.
There are other valuable directions for extension. For example, Callaway and Sant’Anna (2021) and Goodman-Bacon (2021) propose a DID method in a more flexible setting than the previous designs, where there are multiple time points and the treatment may be initiated at any of them. Moreover, De Chaisemartin and D’haultfœuille (2023) develops a DID method when there are more than two groups, and Richardson et al. (2023) proposes the generalized DID to estimate causal effects under alternative assumptions to the parallel trend assumption. While the outcome in the basic DID method is continuous, Tchetgen et al. (2024) deals with the case where the response follows a generalized linear model. Thus, the DID method has been theoretically developed in accordance with real problems, and extensions of the proposed method to these problems are recent challenges. In econometrics, the synthetic control method (Abadie et al. 2010) is a major method for causal inference as well as the DID method, and a method combining it with the DID method has also been developed as synthetic DID (Arkhangelsky et al. 2021). In this method, the objective of the information criterion is what to synthesize and with what weights, and the proposed method cannot necessarily be extended in a similar way, but it is an important topic. Furthermore, propensity score analysis with high-dimensional confounding variables is also a hot topic (e.g., Belloni et al. 2017, Chernozhukov et al. 2018, Ning et al. 2020b), and even in that analysis, robustly estimating the influence of covariates that directly affect the causal effect and selecting covariates that truly affect the causal effect is an essential task to ensure the validity of the analysis. The asymptotic properties of the estimator of causal effects tend to be obtained even when the covariates are high-dimensional as well, and it is surely possible to extend our proposed method for these analyses.

### v3

@@ 5.1 Derivation of model selection criterion
As we did for the case of known propensity scores in Section 4.2, here, we will derive a validated model selection criterion for when the CBD method is used. Let us evaluate
@@
which is the third term of the risk in (16). Substituting $\bm{x}_{i}^{{\rm T}}\bm{\theta}^{*}$ into ${\rm E}[y^{[1]}_{i}-y^{[0]}_{i}\mid\bm{x}_{i},d^{[1]}_{i}=1]$ first and then substituting the right-hand side of (13) into $\hat{\bm{\theta}}^{\rm CBD}-\bm{\theta}^{*}$ yields
$2{\rm E}\bigg[\frac{1}{\sqrt{n}}\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})\{\rho(\bm{t}_{i},\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})\Delta_{i}-\bm{x}_{i}^{{\rm T}}\bm{\theta}^{*}\}\bm{x}_{i}^{{\rm T}}\bigg\{\frac{1}{n}\sum_{j=1}^{n}e^{[1]}(\bm{x}_{j};\hat{\bm{\alpha}}^{\rm CB})\bm{x}_{j}\bm{x}_{j}^{{\rm T}}\bigg\}^{-1}$
@@
Since $n^{-1}\sum_{j=1}^{n}e^{[1]}(\bm{x}_{j};\hat{\bm{\alpha}}^{\rm CB})\bm{x}_{j}\bm{x}_{j}^{{\rm T}}$ converges in probability to
$\bm{L}(\bm{\alpha}^{\dagger})\equiv{\rm E}[e^{[1]}(\bm{x};\bm{\alpha}^{\dagger})\bm{x}\bm{x}^{{\rm T}}],$ (20)
by letting $\bm{u}$ be the Gaussian random vector that is the weak limit of
@@
we can express the weak limit of (16) with expectation removed as
@@
When $\hat{\bm{\theta}}^{\rm CBD}$ is a consistent estimator of $\bm{\theta}^{*}$, it holds that ${\rm E}[\bm{u}\bm{u}^{{\rm T}}]={\rm Var}[\bm{u}]$ since the expectation of $\bm{u}$ becomes $\bm{0}$. Accordingly, the asymptotic evaluation of (16), which is to be used in the information criterion, can be written as
@@
Now let us derive a specific expression for ${\rm Var}[\bm{u}]$. Since the error of the estimator $\hat{\bm{\alpha}}^{\rm CB}$ in the propensity scores is written as (12), we can expand (21) as
@@
[…]
$=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}[e^{[1]}(\bm{x}_{i};\bm{\alpha}^{\dagger})\{\rho(\bm{d}_{i},\bm{x}_{i};\bm{\alpha}^{\dagger})\Delta_{i}-\bm{x}_{i}^{{\rm T}}\bm{\theta}^{*}\}\bm{x}_{i}$
[…]
@@
it can be shown that ${\rm Var}[\bm{u}]=\bm{V}(\bm{\alpha}^{\dagger},\bm{\theta}^{*})$. Thus, the following theorem is obtained.
@@
When the propensity scores are estimated using the CBD method, the third term in (15) is evaluated as
${\rm E}[c^{\rm limit}]=2{\rm tr}\{\bm{L}(\bm{\alpha}^{\dagger})^{-1}\bm{V}(\bm{\alpha}^{\dagger},\bm{\theta}^{*})\},$
where $\bm{L}(\cdot)$ and $\bm{V}(\cdot,\cdot)$ are defined in (20) and (22), respectively.
Since the penalty term in this theorem depends on the true values of the parameters, we can use empirical estimates in practice. Specifically, we will replace $\bm{L}(\bm{\alpha}^{\dagger})$ by $\bm{L}_{n}(\hat{\bm{\alpha}}^{\rm CB})\equiv n^{-1}\sum_{i=1}^{n}\allowbreak e^{[1]}(\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})\bm{x}_{i}\bm{x}_{i}^{{\rm T}}$ and $\bm{V}(\bm{\alpha}^{\dagger},\bm{\theta}^{*})$ by
$\bm{V}_{n}(\hat{\bm{\alpha}}^{\rm CB},\hat{\bm{\theta}}^{\rm CBD})$
$\equiv\frac{1}{n}\sum_{i=1}^{n}[e^{[1]}(\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})\{\rho(\bm{d}_{i},\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})\Delta_{i}-\bm{x}_{i}^{{\rm T}}\hat{\bm{\theta}}^{\rm CBD}\}\bm{x}_{i}$
[…]
where
$\bm{M}_{n}(\bm{\alpha},\bm{\theta})\equiv\frac{1}{n}\sum_{i=1}^{n}\dfrac{\partial}{\partial\bm{\alpha}}e^{[1]}(\bm{x}_{i};\bm{\alpha})\bigg[\frac{d^{[1]}_{i}-1}{\{1-e^{[1]}(\bm{x}_{i};\bm{\alpha})\}^{2}}\Delta_{i}-\bm{x}_{i}^{{\rm T}}\bm{\theta}\bigg]\bm{x}_{i}.$ (23)
As a result, for when the propensity scores are estimated with the CBD method, we propose
$\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})\{\rho(\bm{d}_{i},\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})\Delta_{i}-\bm{x}_{i}^{{\rm T}}\hat{\bm{\theta}}^{\rm CBD}\}^{2}+2{\rm tr}\{\bm{L}_{n}(\hat{\bm{\alpha}}^{\rm CB})^{-1}\bm{V}_{n}(\hat{\bm{\alpha}}^{\rm CB},\hat{\bm{\theta}}^{\rm CBD})\}$ (24)
as the model selection criterion, which is an asymptotically unbiased estimator of the risk in (15). Here, the propensity scores are estimated using covariate balancing, but the information criterion can be derived in the same way when the MLE is used (see Supplementary Material).
@@ 5.2 Numerical experiments
The simulation studies presented in this section examine the accuracy of the approximation for the penalty term and the performance of the model selection criterion in (24) for the proposed CBD method. On the basis of the results in Section 4.4, we will use the identity matrix as the weighting matrix in the generalized method of moments and note that, even when the optimal matrix is used as the weighting matrix, the accuracy of the approximation for the penalty term and the performance of the model selection with comparable to those obtained with the identity matrix (see the Supplementary Material).
Table 4 lists results for data generated as in Cases 1-1 and Cases 1-2 in Section 4.4 to check whether the second term in (24), the penalty term of the proposed criterion, accurately approximates the bias in (16), which we originally wanted to evaluate, for various values of the parameter $\beta^{*}$ and sample size $n$. In all cases, the evaluated penalty terms of the proposed criterion are much closer to the bias than the evaluated penalty terms of the QICw, and its approximation accuracy is high. In contrast, the penalty term of the QICw significantly underestimates the bias in all cases.
Table 4: Bias evaluation for case in which the propensity scores are estimated using the CBD method with the identity matrix as the weighting matrix: True denotes the true value of the bias evaluated by the Monte Carlo method, and Proposal the penalty term of the proposed criterion. The penalty term of the QICw is also listed for reference.
@@
Table 5 lists results for data generated as in Case 2-1, Case 2-2, and Case 2-3 in Section 4.4 and comparing the proposed criterion and the QICw through the forward selection method. The main index of goodness for the model selection criteria is again the empirical estimate of the risk. Unlike Section 4.4, where the propensity scores are known, the proposed criterion always outperforms the QICw in all cases, even in Case 2-2, where relatively few covariates are unnecessary among the candidates. The proposed criterion is especially superior when the candidates include many covariates that do not contribute to the true structure, as in Case 2-1 and Case 2-3 of Section 4.4. The QICw underestimates the penalty term; thus, too many covariates are selected.
Table 5: Comparison of the proposed criterion and the QICw through model selection for case in which the propensity scores are estimated using the CBD method with the identity matrix as the weighting matrix: Risk denotes the empirical estimate of the risk of the selected model, TP the average number of covariates that are true positive, and FP the average number of covariates that are false positive.
@@ 6 Real data analysis
The LaLonde dataset, originally presented in LaLonde (1986), is included in the R package Matching version 4.10.14. In the dataset, the group that took the U.S. job training program in 1976 is denoted as $t=1$ and the group that did not take the program is denoted as $t=0$. The difference in annual income in 1978 after the training for each group is estimated via the ATT. Age (age, years of education (educ), income in 1974 (re74), black (black), hispanic (hisp), married (married), and high school graduate or higher (nodegr) are covariates, i.e. $p=7$, and the outcome variable is the difference between income in 1978 (re78) and income in 1974 (re78). The sample size is $n=445$. The objective of our experiment here is not to extract information from these data, but to determine whether the proposed method makes a significant difference from an existing method. Therefore, to reduce the sample size of the dataset, we divided the data into three parts and analyzed each of them. Specifically, for every integer $k$, the 3k23k-2-th row of data are included in Block 1, the 3k13k-1-th row in Block 2, and the $3k$-th row in Block 3, respectively.
First, we estimate the propensity scores by using the covariate balancing with the identity matrix as the weighting matrix. Then, supposing that the ATT can be written as a linear sum of covariates, we estimate their coefficients by using the SDID approach. In Table 5, the covariates are selected using the proposed criterion or the QICw with the forward selection method and the coefficients of the selected covariates are estimated. The QICw selects all covariates in any block, while the proposed criterion does not select some variables. In particular, the difference between the models selected for Block 1 and Block 3 is rather large. Because we do not know the true structure of the real data, it is impossible to judge which of the two criteria is best; however, we can see that the difference between them is quite large. These results indicates that, in a setting such as here, it is essential to use a model selection criterion that is theoretically valid rather than an intuitive one.
Table 6: Comparison of the proposed criterion and the QICw through model selection for the LaLonde dataset. The numbers with exponential notation are the estimated coefficients for each covariate, and “0” means that the covariate is not selected in the model selection.
@@
QICw 2.38E+4 -1.91E+2 8.22E+2 2.73E-1
3 Proposal -8.50E+2 0E+0 0E+0 0E+0
QICw -8.51E+3 6.13E+0 3.35E+2 2.56E-1
@@ 1 Proposal 0E+0 0E+0 0E+0 0E+0
QICw 3.55E+2 -8.19E+3 -9.41E+2 -1.30E+3
2 Proposal -1.86E+4 -2.34E+4 4.02E+3 -1.26E+4
QICw -1.85E+4 -2.42E+4 4.88E+3 -1.34E+4
3 Proposal 0E+0 -1.03E+4 0E+0 0E+0
QICw 3.84E+3 -6.71E+3 2.92E+3 -5.90E+2
@@ 7 Discussion
Our proposal, the CBD method, is an SDID approach incorporating covariate balancing. In order to ensure double robustness, i.e. to ensure that the ATT is consistently estimated when the model for the propensity score is correctly specified or the model for the changes in the outcomes over time is correctly supposed, we have shown that the second-order moments of the distribution for the covariates should be balanced between the treatment and control groups with the former as the target population. Since covariate balancing usually balances the first-order moments, the value for the fact that the moment is second-order rather than first-order lies more in having been discovered due to its unexpectedness than in how to demonstrate its validity. Moreover, our model selection criterion for the CBD method is an asymptotically unbiased estimator of risk based on the loss function used for the estimation. Specifically, we derived a model selection criterion using just a DID-specific assumption called the conditional parallel trend assumption, i.e., without relying on the ignorable treatment assignment condition, which is often assumed in causal inference. The penalty term has been shown to be considerably different from twice the number of parameters that often appears in the AIC-type model selection criteria. Since no model selection criteria existed for even the basic SDID approach, we developed a model selection criterion not only for the CBD method, but also for when the propensity scores are known or estimated by MLE. Numerical experiments show that the proposed method estimates the ATT more robustly than the method that estimates propensity scores by using MLE and that the proposed criterion clearly reduces the risk targeted by the SDID approach compared with the intuitive generalization of the existing information criterion to the SDID approach. In addition, analyses of real data confirm that there is a large difference between the results of the proposed method and those of the existing method.
Although this paper deals with a simple linear model of covariates as a working model for the outcome, its results can be easily extended to situations where a nonlinear model of covariates should be used, for example, when using the kernel method. In fact, Hazlett (2020) has treated the function $\bm{\phi}(\bm{x}):\mathbb{R}^{p}\rightarrow\mathbb{R}^{q}$ in a reproducing kernel Hilbert space where $q$ may be larger than $p$ or $n$. They assumed that the outcome could be represented by a linear model of $\bm{\phi}(\bm{x})$, which in our setting can be written as
${\rm E}[y^{[1]}(1)-y^{[0]}(1)\mid\bm{x},d^{[1]}=1]=\bm{\phi}(\bm{x})^{{\rm T}}\bm{\theta}.$ (25)
@@
$\sum_{i=1}^{n}\bigg\{\frac{d^{[1]}_{i}\bm{\phi}(\bm{x}_{i})}{e^{[1]}(\bm{x}_{i};\bm{\alpha})}-\frac{d^{[0]}_{i}\bm{\phi}(\bm{x}_{i})}{e^{[0]}(\bm{x}_{i};\bm{\alpha})}\bigg\}=\bm{0}$
as the moment conditions for estimating the propensity scores, they propose the kernel balancing method, which can flexibly estimate causal effects even if the relationship between the covariate and the outcome is nonlinear. For making a doubly robust estimation using the kernel balancing in the setting of this paper, together with the assumption in (25), let us additionally suppose that the change in the outcomes follows
${\rm E}[\Delta^{[k]}\mid\bm{x},d^{[k]}=1]=\bm{\phi}(\bm{x})^{{\rm T}}\bm{\beta}^{[k]*}.$
Let $k(\cdot,\cdot):\mathbb{R}^{p}\times\mathbb{R}^{p}\rightarrow\mathbb{R}$ be the kernel function corresponding to $\bm{\phi}(\bm{x})$, and let $\bm{K}$ be a $p\times p$ semi-definite symmetric matrix whose $(l,m)$-component is $k(x_{l},x_{m})$. If we estimate the parameter $\bm{\alpha}$ that satisfies
@@
and then estimate $\bm{\theta}$ and derive a model selection criterion based on it, it would be a natural extension of the proposed method.
There are other valuable directions of study. For example, Callaway and Sant’Anna (2021) and Goodman-Bacon (2021) proposed a DID approach in a more flexible setting than the original one, where there are multiple time points and treatment can be initiated at any of them. Moreover, De Chaisemartin and D’haultfœuille (2023) developed a DID approach for when there are more than two groups, and Richardson et al. (2023) proposed one to estimate causal effects under alternative assumptions to the parallel trend assumption. While the outcome in the basic DID approach is continuous, Tchetgen et al. (2024) dealt with the case where the outcome follows a generalized linear model. In this way, the DID approach has been theoretically developed in accordance with actual problems, so we believe that extending the proposed method to enable it to deal with those problems would thus be a potentially profitable goal for the near future. In econometrics, synthetic control (Abadie et al. 2010) has also become standard in causal inference, and a combination of the two has been developed as synthetic DID (Arkhangelsky et al. 2021). Here, because the objective of the information criterion is what to synthesize and with what weights, the proposed method cannot necessarily be used; however, an investigation along the lines presented in this paper might prove fruitful. Furthermore, propensity score analysis with high-dimensional covariates is also an important topic (e.g., Belloni et al. 2017, Chernozhukov et al. 2018, Ning et al. 2020b), wherein selecting covariates that truly have impact on causal effects and robustly estimating these effects is an essential task to ensure the validity of the analysis. Even when the covariates are high-dimensional, asymptotic properties of the estimator of causal effects tend to be obtained as well, so it is surely possible to extend our method.
Disclosure statement
No potential conflict of interest was reported by the authors.
Data availability statement
The LaLonde dataset is available in the R package Matching version 4.10.14 (Sekhon 2011). Example R scripts used for the simulations and real data analyses are available from the authors.
Funding
Yoshiyuki Ninomiya gratefully acknowledges support from JSPS KAKENHI (grant numbers 23H00809 and 23K18471)